# neurips_2025.bib

### v1

@@ D.3 Action Adapter
if training then
       Remove self-match: keep indices $[1:K]$ ;
else
       Use top $K$ except last: $[0:K]$ ;
$\mathbf{W}\leftarrow\texttt{Softmax}(\texttt{topk\_sim}/T)$ ;
@@
$\hat{\mathbf{A}}_{t}\leftarrow\sum\nolimits_{k=1}^{K}\mathbf{W}_{k}\cdot\mathbf{A}_{\text{top},k}$ ;
if $\max(\texttt{topk\_sim})<B$ then
       $\hat{\mathbf{A}}_{t}\leftarrow\mathbf{0}$ ;
        // Mask if too dissimilar
return $\hat{\mathbf{A}}_{t}$ ;
Algorithm 1 Action Adaptation: Similarity-based Action Retrieval
Appendix E Details of Real-world Experiments
@@ E.1 Experimental Setup
Model deployment. We evaluate our method on several real-world tasks using the Franka Research 3 Robot equipped with a 7-DoF arm and a 1-DoF Panda Hand gripper. Visual input is provided by a single Intel D435i camera configured in an eye-on-base setup, as shown in Figure 13. Both OpenVLA [openvla] and CronusVLA require over 10 GB of memory in bfloat16 precision; thus, all models are deployed on an online A100 GPU server, while a local machine handles robot control. Although our method and DP3 support higher frequencies, we limit both data collection and action execution to 5 Hz to match OpenVLA’s speed constraints and mitigate communication latency. Finetuning data are collected via human teleoperation using a SpaceMouse device.
@@
Training details. For CronusVLA, we follow the finetuning protocol from the LIBERO experiments, initializing CronusVLA-7B with post-trained weights and conducting cross-embodiment finetuning on our collected demonstrations. For DP3[dp3], we adopt the updated instruction-conditioned architecture from[ke20243d] and implement it ourselves. We report the best-performing checkpoint of DP3. For OpenVLA [openvla], we follow the official training protocol, fully finetuning the model until the token accuracy reaches 95%, and report its best checkpoint. All models are trained using a consistent data augmentation strategy.
E.2 Tasks Settings
@@
Simple Pick-and-Place. These tasks involve straightforward pick-and-place actions with at most two sub-tasks. Detailed success rates are shown in Table 11, visualizations are shown in Figure 14 and our video demo. We evaluate the model’s learning capability and spatial generalization through simple manipulation tasks:
@@
  • Pick objects. This task assesses the model’s fundamental abilities in localization, grasping, and spatial generalization. Four target objects are used: an eggplant, a carrot, a wooden cube, and a bowl. The eggplant and carrot are relatively easy to grasp, while the cube and bowl demand higher placement precision to prevent arm jamming or slippage. Object positions are changed within a region centered on the plate. Illustrated in Figure 14 (a).
@@
  • Put the carrot on the plate. This task primarily evaluates grasping precision and placement accuracy for the single-object pick-and-place. The robot must grasp a carrot from a tabletop region on the right, where objects are positioned with varying locations and orientations, and place it into the designated bowl. Illustrated in Figure 14 (b).
@@
  • Stack the red cube on the blue cube. This task poses a greater challenge than the previous pick-and-place setup. The robot is required to accurately grasp a red cube from a designated area and stack it on a randomly positioned blue cube. The varying positions of both cubes across trials increase task complexity, especially for models relying exclusively on third-person visual observations. Illustrated in Figure 14 (c).
@@
  • Stack the red cup into the green one. This task evaluates the model’s capabilities of instruction following and spatial perception. We introduce considerable spatial variability by randomly placing the red, green, and yellow cups within three predefined regions and frequently permuting their positions. During data collection, either the red or yellow cup is randomly placed into the green cup, with corresponding language instructions annotated. And evaluation is restricted to the instruction, "stack the red cup into the green one". Illustrated in Figure 14 (d).
Long-horizon Tasks. We evaluate tasks involving multi-step object manipulation with at least two sub-tasks to evaluate the model’s proficiency in sequential execution and long-horizon task composition. Detailed success rates are shown in Table 12, visualizations are shown in Figure 15 and our video demo:
  • Put the multiple objects on the plate in order. This task assesses the model’s ability of multi-step execution and instructions following. It extends the single-object pick-and-place setup to a multi-object setting. During data collection, one object is randomly selected for each grasping step until all objects are relocated, with varying relative placements in the bowl. In evaluation, the instruction of grasping order is fixed. Illustrated in Figure 15 (a).
  • Open and then close the drawer. This task evaluates the model’s ability to manipulate articulated objects through sequential actions, emphasizing precise localization of interaction points. The robot must grasp a drawer handle, pull to open, release and reposition the gripper, and push to close the drawer. During both training and evaluation, the drawer is rotated differently, with its initial opening set to one of three predefined states. Success is defined by the correct execution of both opening and closing. Data collection is split into two stages, while evaluation follows a unified sequential protocol. Illustrated in Figure 15 (b).
  • Open the drawer and place the carrot into the drawer. This task assesses the model’s stability across a wide spatial range and its capability in long-horizon manipulation. The robot must sequentially pull open a drawer, grasp an object from the plate, and place it into the drawer. Variations include the carrot’s position and orientation within the plate, as well as the drawer’s initial opening state. Illustrated in Figure 15 (c).
  • Press the buttons in order. This evaluation primarily measures the model’s ability to disambiguate observations and accurately follow instructions. The robot sequentially presses three color-coded buttons whose positions vary within a defined range. The pressing order of testing uses a fixed sequence: red, yellow, then green. The task introduces substantial visual ambiguity, as similar observations occur at different stages, such as before and after button presses. Illustrated in Figure 15 (d).
Generalization and Robustness. We evaluate the generalization and robustness of different policies. All policies are evaluated on the Put the carrot on the plate task and trained on the data collected in all seven tasks of Simple Pick-and-place. Detailed success rates are shown in Table 13, visualizations are shown in Figure 16 and our video demo:
  • Material variance. This experiment aims to evaluate the model’s generalization across objects. Carrots differing in appearance and material properties are used, as is shown in Figure 16 (a). We conduct the evaluation using the same spatial variance settings as in the Put the carrot on the plate task.
  • Human interference. This experiment assesses the model’s robustness under manual disturbances. In each test trial, irrelevant or similar objects are thrown to introduce interference, as shown in Figure 16 (b).
  • Camera occlusion. This experiment evaluates the model’s robustness to incomplete observations. The camera view is occluded at a fixed frequency to simulate partial or corrupted input, as illustrated in Figure 16 (c).
  • Unseen objects. This experiment assesses the model’s generalization to grasping different objects. We test objects not seen during training, placed in similar varying positions, as shown in Figure 16 (d).
  • Unseen background. This experiment evaluates the model’s robustness to varying backgrounds. The background is changed from a plain white surface to a textured one, as shown in Figure 16 (e).
  • Unseen distractors. This experiment assesses the impact of distractor objects on the model’s performance. A large number of unrelated additional objects, such as toys, are placed on the table, as shown in Figure 16 (f).
  • Variable lighting. This experiment assesses the model’s robustness to varying lighting conditions by employing periodic flashing lights, as shown in Figure 16 (g).
@@
Ours 72.0 84.0 64.0 40.0 80.0 80.0 56.0 48.0 28.0 16.0
@@
Ours 52.0 28.0 20.0 72.0 64.0 72.0 60.0 96.0 92.0 88.0
@@
Appendix F Failure Cases
@@
Despite the strong performance of CronusVLA in both simulation benchmarks and real-world experiments, notable failure cases remain, highlighting opportunities for further improvement. As illustrated in Figure 17, we present four representative examples. In (A), a real-world stack cups task, the model reaches a graspable position but makes unnecessary adjustments. This may result from the limitations of a single third-person view, particularly in cluttered scenes where occlusions and viewpoint variance impair perception. Incorporating more cluttered-scene data and adding wrist-view input could improve robustness. In (B), the model struggles with depth estimation when opening distant drawers; despite multiple attempts, the model consistently fails to accurately grasp the bottom drawer. This underscores the importance of precise depth perception in such tasks. In (C), although CronusVLA selects the correct action, the robot arm knocks the target object (an apple) away due to occlusion-induced collisions. We attribute the failure to the model’s lack of explicit obstacle avoidance under occlusion. In (D), certain behaviors acceptable in the SimplerEnv simulation result in unrealistic physical outcomes, such as a slight touch causing a can to flip and roll excessively. This suggests a real-to-sim gap that needs further investigation.
Appendix G Complexity Analysis
@@
Appendix H Limitation and Future Works
Efficient utilization of inter-frame information. Inspired by the effectiveness of multi-frame inputs in low-level action prediction, we introduce a feature-level multi-frame modeling approach that utilizes motion features extracted by the VLA module. However, in scenarios where observations and instructions remain largely unchanged across frames, large language models (LLMs) tend to process redundant tokens, leading to considerable spatial redundancy. Mitigating this redundancy by capturing inter-frame differences can enhance model efficiency and lower training costs. A promising direction involves exploring cache-like mechanisms that leverage inter-frame image differences and the compositional nature of LLMs to improve efficiency further.
@@ Appendix H Limitation and Future Works
Appendix I Broader Impact
@@
Appendix J More Backgrounds
@@
Vision-language data. Our 7B model is based on the Prismatic [prismatic] backbone, which combines the LLaMA2 [llama2] language model with DINOv2 [dinov2] and SigLIP [siglip] as vision encoders. LLaMA2 is pretrained on approximately 2 trillion language tokens, while DINOv2 and SigLIP are trained on large-scale visual datasets. Prismatic is trained for vision-language alignment using datasets such as LAION [schuhmann2021laion], Conceptual Captions [sharma2018conceptual], and SBU Captions [ordonez2011im2text], totaling roughly 558K image-caption pairs. It is further instruction-tuned on 665K multimodal data, including LLaVA Synthetic Data [llava], VQA datasets [goyal2017making, hudson2019gqa], referring expression datasets [kazemzadeh2014referitgame, krishna2017visual], and so on.
@@
Vision-language-action data. For VLA data, we follow the pretraining protocols of Octo [octo] and OpenVLA [openvla], utilizing 27 datasets from Open X-Embodiment [oxe]. These datasets primarily comprise single-arm demonstrations with at least one third-person observation and corresponding end-effector actions. Note that some subsets, such as Kuka [kalashnikov2018scalable], may contain imperfect language instructions. And as shown in [cogact, openvla], Droid [droid] and Language Table [lynch2023interactive] datasets may lead to a potential out-of-distribution issue. Especially, the Fractal dataset is a large-scale collection of open-world manipulation demonstrations, comprising approximately 87k episodes and 3.8M images. These demonstrations were collected using the Google Robot platform, providing a diverse set of real-world robotic interactions. Bridge-v2 is another extensive dataset designed to facilitate scalable robot learning. It contains 60k trajectories and 2.1M images collected across different environments using the WidowX Robot platform. All data is stored in the RLDS format [ramos2021rlds] and mixed using predefined ratios during training.
@@
Training for discrete action. The cross-entropy loss function is defined as: $\mathcal{L}_{\text{CE}}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid y_{<t},x)$, where $y_{t}$ represents the target token at position $t$, $y_{<t}$ denotes the sequence of preceding tokens, $x$ encapsulates the input modalities (including images and language instructions), and $P_{\theta}$ is the model’s predicted probability distribution parameterized by $\theta$. For action prediction, we follow RT-2 [RT2] and OpenVLA [openvla], employing a discretization strategy wherein each dimension of the robot’s continuous action space is divided into 256 uniform bins. This transforms continuous action values into discrete tokens, facilitating their integration into the language modeling framework. To accommodate these action tokens within the model’s vocabulary, OpenVLA repurposes the 256 least frequently used tokens in the Llama tokenizer’s vocabulary, replacing them with the newly introduced action tokens.
@@
SimplerEnv. In the simulation, we conduct the majority of our experiments within SimplerEnv [simplerenv], a benchmark designed to evaluate the capabilities of models in performing various tasks with the WidowX Robot and the Google Robot environment, which can effectively evaluation the cross-embodiment generalization and exhibit a strong performance correlation between simulator and real-world scenarios. For the Google Robot environment, SimplerEnv includes two experimental settings. Visual Matching (VM) strictly follows a real-to-sim replication, assessing the model’s ability to learn demonstration trajectories while testing its spatial generalization through viewpoint and position variations. Variant Aggregation (VA) introduces environmental variations in background, lighting, distractors, and textures, evaluating the policy’s adaptability and robustness in a zero-shot manner. Both of them primarily include tasks such as picking a coke can, moving near, opening/closing the drawer, opening the drawer, and placing the apple in. For the WidowX Robot environment, only the Visual Matching (VM) setting is used, ensuring a faster evaluation aligned with Bridge-v2 [bridgev2]. It mainly includes tasks such as putting the carrot on the plate, putting the spoon on the towel, stacking the green block on the yellow block, and putting the eggplant in the yellow basket. We evaluate the SimplerEnv following the setting in  [cogact].
@@
LIBERO. It comprises four task suites, including LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long, to assess specific aspects of knowledge transfer in lifelong robot learning. For LIBERO-Spatial, it focuses on spatial reasoning, involving manipulating identical objects where the primary challenge lies in discerning spatial relationships. LIBERO-Object emphasis is on the object-centric manipulation, requiring the robot to interact with unique objects and highlighting the transfer of object-specific knowledge. LIBERO-Goal evaluates goal-conditioned adaptability, requiring the robot to adjust its actions to achieve different goals, testing its understanding of task objectives. LIBERO-Long is designed to challenge long-term planning and execution, comprising multi-stage tasks that require sustained attention and coordination. Each task suite comprises 10 tasks with 50 human-teleoperated demonstrations. We follow the data cleaning procedure of  [openvla], which filters out corrupted samples and standardizes the data format to RLDS.

### v2

@@ E.3 Efficiency Analysis
Inference
@@
Inference
[…]
Inference
[…]
Training
Step (K)
OpenVLA 1 24.7 5.2 192.3 15.0 295
+MultiFrame 7 32.4 3.1 263.2 17.2 50
SpatialVLA 1 51.2 2.4 322.5 8.6 200
Ours-0.5B 4 56.0 11.1 90.1 3.7 40
Ours-7B 7 70.9 8.7 114.9 15.1 50
Table 18: Comparison of performance, efficiency, and cost. Inference is conducted on A100 GPUs without chunking, and success rates on SimplerEnv are reported. Training steps are summarized from official setups. Different VLA models with varying frame numbers and latency–accuracy trade-off are also listed.
Appendix F Details of Real-world Experiments
@@ F.1 Experimental Setup
Model deployment. We evaluate our method on several real-world tasks using the Franka Research 3 Robot equipped with a 7-DoF arm and a 1-DoF Panda Hand gripper. Visual input is provided by a single Intel D435i camera configured in an eye-on-base setup, as shown in Fig.16. Both OpenVLA (Kim et al. 2024) and CronusVLA require over 10 GB of memory in bfloat16 precision; thus, all models are deployed on an online A100 GPU server, while a local machine handles robot control. Although our method and DP3 support higher frequencies, we limit both data collection and action execution to 5 Hz to match OpenVLA’s speed constraints and mitigate communication latency. Finetuning data are collected via human teleoperation using a SpaceMouse device.
@@
Training details. For CronusVLA, we follow the finetuning protocol from the LIBERO experiments, initializing CronusVLA-7B with post-trained weights and conducting cross-embodiment finetuning on our collected demonstrations. For DP3(Ze et al. 2024), we adopt the updated instruction-conditioned architecture from(Ke, Gkanatsios, and Fragkiadaki 2024) and implement it ourselves. We report the best-performing checkpoint of DP3. For OpenVLA (Kim et al. 2024), we follow the official training protocol, fully finetuning the model until the token accuracy reaches 95%, and report its best checkpoint. All models are trained using a consistent data augmentation strategy.
F.2 Tasks Settings
@@
Simple Pick-and-Place. These tasks involve straightforward pick-and-place actions with at most two sub-tasks. Detailed success rates are shown in Tab.19, visualizations are shown in Fig.17 and our video demo. We evaluate the model’s learning capability and spatial generalization through simple manipulation tasks:
@@
  • Pick objects. This task assesses the model’s fundamental abilities in localization, grasping, and spatial generalization. Four target objects are used: an eggplant, a carrot, a wooden cube, and a bowl. The eggplant and carrot are relatively easy to grasp, while the cube and bowl demand higher placement precision to prevent arm jamming or slippage. Object positions are changed within a region centered on the plate. Illustrated in Fig.17 (a).
@@
  • Put the carrot on the plate. This task primarily evaluates grasping precision and placement accuracy for the single-object pick-and-place. The robot must grasp a carrot from a tabletop region on the right, where objects are positioned with varying locations and orientations, and place it into the designated bowl. Illustrated in Fig.17 (b).
@@
  • Stack the red cube on the blue cube. This task poses a greater challenge than the previous pick-and-place setup. The robot is required to accurately grasp a red cube from a designated area and stack it on a randomly positioned blue cube. The varying positions of both cubes across trials increase task complexity, especially for models relying exclusively on third-person visual observations. Illustrated in Fig.17 (c).
@@
  • Stack the red cup into the green one. This task evaluates the model’s capabilities of instruction following and spatial perception. We introduce considerable spatial variability by randomly placing the red, green, and yellow cups within three predefined regions and frequently permuting their positions. During data collection, either the red or yellow cup is randomly placed into the green cup, with corresponding language instructions annotated. And evaluation is restricted to the instruction, ”stack the red cup into the green one”. Illustrated in Fig.17 (d).
@@
[…]
Long-horizon Tasks. We evaluate tasks involving multi-step object manipulation with at least two sub-tasks to evaluate the model’s proficiency in sequential execution and long-horizon task composition. Detailed success rates are shown in Tab.20, visualizations are shown in Fig.18 and our video demo:
  • Put the multiple objects on the plate in order. This task assesses the model’s ability of multi-step execution and instructions following. It extends the single-object pick-and-place setup to a multi-object setting. During data collection, one object is randomly selected for each grasping step until all objects are relocated, with varying relative placements in the bowl. In evaluation, the instruction of grasping order is fixed. Illustrated in Fig.18 (a).
  • Open and then close the drawer. This task evaluates the model’s ability to manipulate articulated objects through sequential actions, emphasizing precise localization of interaction points. The robot must grasp a drawer handle, pull to open, release and reposition the gripper, and push to close the drawer. During both training and evaluation, the drawer is rotated differently, with its initial opening set to one of three predefined states. Success is defined by the correct execution of both opening and closing. Data collection is split into two stages, while evaluation follows a unified sequential protocol. Illustrated in Fig.18 (b).
  • Open the drawer and place the carrot into the drawer. This task assesses the model’s stability across a wide spatial range and its capability in long-horizon manipulation. The robot must sequentially pull open a drawer, grasp an object from the plate, and place it into the drawer. Variations include the carrot’s position and orientation within the plate, as well as the drawer’s initial opening state. Illustrated in Fig.18 (c).
  • Press the buttons in order. This evaluation primarily measures the model’s ability to disambiguate observations and accurately follow instructions. The robot sequentially presses three color-coded buttons whose positions vary within a defined range. The pressing order of testing uses a fixed sequence: red, yellow, then green. The task introduces substantial visual ambiguity, as similar observations occur at different stages, such as before and after button presses. Illustrated in Fig.18 (d).
@@
[…]
Generalization and Robustness. We evaluate the generalization and robustness of different policies. All policies are evaluated on the Put the carrot on the plate task and trained on the data collected in all seven tasks of Simple Pick-and-place. Detailed success rates are shown in Tab.21, visualizations are shown in Fig.19 and our video demo:
  • Material variance. This experiment aims to evaluate the model’s generalization across objects. Carrots differing in appearance and material properties are used, as is shown in Fig.19 (a). We conduct the evaluation using the same spatial variance settings as in the Put the carrot on the plate task.
  • Human interference. This experiment assesses the model’s robustness under manual disturbances. In each test trial, irrelevant or similar objects are thrown to introduce interference, as shown in Fig.19 (b).
  • Camera occlusion. This experiment evaluates the model’s robustness to incomplete observations. The camera view is occluded at a fixed frequency to simulate partial or corrupted input, as illustrated in Fig.19 (c).
  • Unseen objects. This experiment assesses the model’s generalization to grasping different objects. We test objects not seen during training, placed in similar varying positions, as shown in Fig.19 (d).
  • Unseen background. This experiment evaluates the model’s robustness to varying backgrounds. The background is changed from a plain white surface to a textured one, as shown in Fig.19 (e).
  • Unseen distractors. This experiment assesses the impact of distractor objects on the model’s performance. A large number of unrelated additional objects, such as toys, are placed on the table, as shown in Fig.19 (f).
  • Variable lighting. This experiment assesses the model’s robustness to varying lighting conditions by employing periodic flashing lights, as shown in Fig.19 (g).
@@
Appendix G Failure Cases
@@
Despite the strong performance of CronusVLA in both simulation benchmarks and real-world experiments, notable failure cases remain, highlighting opportunities for further improvement. As illustrated in Fig.20, we present four representative examples. In (A), a real-world stack cups task, the model reaches a graspable position but makes unnecessary adjustments. This may result from the limitations of a single third-person view, particularly in cluttered scenes where occlusions and viewpoint variance impair perception. Incorporating more cluttered-scene data and adding wrist-view input could improve robustness. In (B), the model struggles with depth estimation when opening distant drawers; despite multiple attempts, the model consistently fails to accurately grasp the bottom drawer. This underscores the importance of precise depth perception in such tasks. In (C), although CronusVLA selects the correct action, the robot arm knocks the target object (an apple) away due to occlusion-induced collisions. We attribute the failure to the model’s lack of explicit obstacle avoidance under occlusion. In (D), certain behaviors acceptable in the SimplerEnv simulation result in unrealistic physical outcomes, such as a slight touch causing a can to flip and roll excessively. This suggests a real-to-sim gap that needs further investigation.
Appendix H Complexity Analysis
@@
Appendix I Limitation and Future Works
Efficient utilization of inter-frame information. Inspired by the effectiveness of multi-frame inputs in low-level action prediction, we introduce a feature-level multi-frame modeling approach that utilizes learnable features extracted by the VLA module. However, in scenarios where observations and instructions remain largely unchanged across frames, large language models (LLMs) tend to process redundant tokens, leading to considerable spatial redundancy. Mitigating this redundancy by capturing inter-frame differences can enhance model efficiency and lower training costs. A promising direction involves exploring cache-like mechanisms that leverage inter-frame image differences and the compositional nature of LLMs to improve efficiency further.
@@ Appendix I Limitation and Future Works
Appendix J Broader Impact
@@
Appendix K More Backgrounds
@@
Vision-language data. Our 7B model is based on the Prismatic (Karamcheti et al. 2024) backbone, which combines the LLaMA2 (Touvron et al. 2023) language model with DINOv2 (Oquab et al. 2023) and SigLIP (Zhai et al. 2023) as vision encoders. LLaMA2 is pretrained on approximately 2 trillion language tokens, while DINOv2 and SigLIP are trained on large-scale visual datasets. Prismatic is trained for vision-language alignment using datasets such as LAION (Schuhmann et al. 2021), Conceptual Captions (Sharma et al. 2018), and SBU Captions (Ordonez, Kulkarni, and Berg 2011), totaling roughly 558K image-caption pairs. It is further instruction-tuned on 665K multimodal data, including LLaVA Synthetic Data (Liu et al. 2023), VQA datasets (Goyal et al. 2017; Hudson and Manning 2019), referring expression datasets (Kazemzadeh et al. 2014; Krishna et al. 2017), and so on.
@@
Vision-language-action data. For VLA data, we follow the pretraining protocols of Octo (Octo Model Team et al. 2024) and OpenVLA (Kim et al. 2024), utilizing 27 datasets from Open X-Embodiment (O’Neill et al. 2024). These datasets primarily comprise single-arm demonstrations with at least one third-person observation and corresponding end-effector actions. Note that some subsets, such as Kuka (Kalashnikov et al. 2018), may contain imperfect language instructions. And as shown in (Li et al. 2024a; Kim et al. 2024), Droid (Khazatsky et al. 2024) and Language Table (Lynch et al. 2023) datasets may lead to a potential out-of-distribution issue. Especially, the Fractal dataset is a large-scale collection of open-world manipulation demonstrations, comprising approximately 87k episodes and 3.8M images. These demonstrations were collected using the Google Robot platform, providing a diverse set of real-world robotic interactions. Bridge-v2 is another extensive dataset designed to facilitate scalable robot learning. It contains 60k trajectories and 2.1M images collected across different environments using the WidowX Robot platform. All data is stored in the RLDS format (Ramos et al. 2021) and mixed using predefined ratios during training.
@@
Training for discrete action. The cross-entropy loss function is defined as: $\mathcal{L}_{\text{CE}}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid y_{<t},x)$, where $y_{t}$ represents the target token at position $t$, $y_{<t}$ denotes the sequence of preceding tokens, $x$ encapsulates the input modalities (including images and language instructions), and $P_{\theta}$ is the model’s predicted probability distribution parameterized by $\theta$. For action prediction, we follow RT-2 (Brohan et al. 2023) and OpenVLA (Kim et al. 2024), employing a discretization strategy wherein each dimension of the robot’s continuous action space is divided into 256 uniform bins. This transforms continuous action values into discrete tokens, facilitating their integration into the language modeling framework. To accommodate these action tokens within the model’s vocabulary, OpenVLA repurposes the 256 least frequently used tokens in the Llama tokenizer’s vocabulary, replacing them with the newly introduced action tokens.
@@
SimplerEnv. In the simulation, we conduct the majority of our experiments within SimplerEnv (Li et al. 2024b), a benchmark designed to evaluate the capabilities of models in performing various tasks with the WidowX Robot and the Google Robot environment, which can effectively evaluation the cross-embodiment generalization and exhibit a strong performance correlation between simulator and real-world scenarios. For the Google Robot environment, SimplerEnv includes two experimental settings. Visual Matching (VM) strictly follows a real-to-sim replication, assessing the model’s ability to learn demonstration trajectories while testing its spatial generalization through viewpoint and position variations. Variant Aggregation (VA) introduces environmental variations in background, lighting, distractors, and textures, evaluating the policy’s adaptability and robustness in a zero-shot manner. Both of them primarily include tasks such as picking a coke can, moving near, opening/closing the drawer, opening the drawer, and placing the apple in. For the WidowX Robot environment, only the Visual Matching (VM) setting is used, ensuring a faster evaluation aligned with Bridge-v2 (Walke et al. 2023). It mainly includes tasks such as putting the carrot on the plate, putting the spoon on the towel, stacking the green block on the yellow block, and putting the eggplant in the yellow basket. We evaluate the SimplerEnv following the setting in  (Li et al. 2024a).
@@
LIBERO. It comprises four task suites, including LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long, to assess specific aspects of knowledge transfer in lifelong robot learning. For LIBERO-Spatial, it focuses on spatial reasoning, involving manipulating identical objects where the primary challenge lies in discerning spatial relationships. LIBERO-Object emphasis is on the object-centric manipulation, requiring the robot to interact with unique objects and highlighting the transfer of object-specific knowledge. LIBERO-Goal evaluates goal-conditioned adaptability, requiring the robot to adjust its actions to achieve different goals, testing its understanding of task objectives. LIBERO-Long is designed to challenge long-term planning and execution, comprising multi-stage tasks that require sustained attention and coordination. Each task suite comprises 10 tasks with 50 human-teleoperated demonstrations. We follow the data cleaning procedure of  (Kim et al. 2024), which filters out corrupted samples and standardizes the data format to RLDS.